40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Nov 24 22:14:08 2012, Koji, Update, General, AP table cleaning up APTable121124.pdfAPTable121124.graffle.zip
    Reply  Tue Nov 27 03:26:00 2012, Jenne, Update, General, POP beam on POP55, POP camera 
    Reply  Tue Nov 27 17:45:42 2012, Steve, Update, General, spare optics of AP table moved to cabinet IMG_1819.JPG
    Reply  Wed Mar 28 16:44:27 2018, Steve, Update, General, AP table today AP_Table_20180328.png
Message ID: 7740     Entry time: Sat Nov 24 22:14:08 2012     Reply to this: 7752   7755   13713
Author: Koji 
Type: Update 
Category: General 
Subject: AP table cleaning up 

On Wednesday (21st) night, I checked the AP table as I wanted to try PRMI locking. 
It was difficult to work with the table as there were so many unnecessary components on it.
Also the beams went through complicated paths as they have funny angles. 

So I decided to clean up of IMC REFl WFS, IFO REFL, and IFO AS paths.
I found that the AS beam was highly astigmatic as the beam went through a (too-much-) tilted lens.
I made several blocked optical paths for REFL and AS for future extension of the detection system.

The current status of the table was uploaded below.

The optical spectrum analyzers and the aux NPRO were left untouched but they should be moved
somewhere (either on the table or outside) which does not disturb the other optical paths.
 


After the cleaning, I started locking PRMI. I could lock PRMI stably. But I could not figure out how
the intra-cavity mode looked like as I did not have the POP camera. The power recycling
gain was not quantitatively evaluated as I did not have POP and I wasn't sure how the beam was aligned at POX/POY.


We need to know:

- Quantitative evaluation of the beam shape in the PRC

- Quantitative evaluation of the power recycling gain

Some obvious things to be fixed

- The POX whitening filters seem not switching. This issue should be checked at the circuit module itself and at the BIO.

- The POX beam is not well focused on the PD. This was particularly clear when PRMI was locked with carrier.

- The POP beam is going nowhere. We need POP55 and POP CCD for diagnoses.

I haven't checked ITMY table.

Attachment 1: APTable121124.pdf  6.390 MB  | Hide | Hide all
APTable121124.pdf
Attachment 2: APTable121124.graffle.zip  6.370 MB
ELOG V3.1.3-